$SPAD/src/input unit-macro.input

Timothy Daly December 3, 2016

Abstract

Test modifications to the $pfMacros internal state variable

Contents

2

— * —

)set break break
)sys rm -f unit-macro.output
)spool unit-macro.output
)set message test on
)set message auto off
)clear all
--S 1 of 9
)lisp (trace |clearParserMacro| |displayMacro| |getParserMacroNames|)
--R
--RValue = (|clearParserMacro| |displayMacro| |getParserMacroNames|)
--E 1
--S 2 of 9
a ==> 3
--R
--R                                                                   Type: Void
--E 2
--S 3 of 9
a ==> 4
--R
--R                                                                   Type: Void
--E 3
--S 4 of 9
b ==> 7
--R
--R                                                                   Type: Void
--E 4
--S 5 of 9
)d macros
--R
--R  1> (|getParserMacroNames|)
--R  <1 (|getParserMacroNames| (|b| |a|))
--R
--RUser-defined macros:
--R   b ==> 7
--R   a ==> 4
--R
--RSystem-defined macros:
--R  1> (|displayMacro| |%e|)
--R   macro %e () == exp(1)
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%i|)
--R   macro %i () == complex(0,1)

--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%infinity|)
--R   macro %infinity () == infinity()
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%minusInfinity|)
--R   macro %minusInfinity () == minusInfinity()
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%pi|)
--R   macro %pi () == pi()
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%plusInfinity|)
--R   macro %plusInfinity () == plusInfinity()
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| SF)
--R   macro SF () == DoubleFloat()
--R  <1 (|displayMacro| NIL)
--E 5
--S 6 of 9
)clear prop a
--R
--R  1> (|getParserMacroNames|)
--R  <1 (|getParserMacroNames| (|b| |a|))
--R  1> (|clearParserMacro| |a|)
--I  <1 (|clearParserMacro| ((|b| |mbody| ((|integer| (|posn| ...
--E 6
--S 7 of 9
)d macros
--R
--R  1> (|getParserMacroNames|)
--R  <1 (|getParserMacroNames| (|b| |a|))
--R
--RUser-defined macros:
--R   b ==> 7
--R   a ==> 3
--R
--RSystem-defined macros:
--R  1> (|displayMacro| |%e|)
--R   macro %e () == exp(1)
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%i|)
--R   macro %i () == complex(0,1)
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%infinity|)
--R   macro %infinity () == infinity()
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%minusInfinity|)
--R   macro %minusInfinity () == minusInfinity()
--R  <1 (|displayMacro| NIL)

--R  1> (|displayMacro| |%pi|)
--R   macro %pi () == pi()
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%plusInfinity|)
--R   macro %plusInfinity () == plusInfinity()
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| SF)
--R   macro SF () == DoubleFloat()
--R  <1 (|displayMacro| NIL)
--E 7
--S 8 of 9
)clear prop a
--R
--R  1> (|getParserMacroNames|)
--R  <1 (|getParserMacroNames| (|b| |a|))
--R  1> (|clearParserMacro| |a|)
--I  <1 (|clearParserMacro| ((|b| |mbody| ((|integer| ...
--E 8
--S 9 of 9
)d macros
--R
--R  1> (|getParserMacroNames|)
--R  <1 (|getParserMacroNames| (|b|))
--R
--RUser-defined macros:
--R   b ==> 7
--R
--RSystem-defined macros:
--R  1> (|displayMacro| |%e|)
--R   macro %e () == exp(1)
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%i|)
--R   macro %i () == complex(0,1)
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%infinity|)
--R   macro %infinity () == infinity()
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%minusInfinity|)
--R   macro %minusInfinity () == minusInfinity()
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%pi|)
--R   macro %pi () == pi()
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| |%plusInfinity|)
--R   macro %plusInfinity () == plusInfinity()
--R  <1 (|displayMacro| NIL)
--R  1> (|displayMacro| SF)
--R   macro SF () == DoubleFloat()

--R  <1 (|displayMacro| NIL)
--E 9
)spool
)lisp (bye)

———-

References^{[1] nothing}

7